An inventor claims to have invented a machine that is capable of manufacturing 50 jg / min of ice at -4 °C, from water at 20 °C and 1 bar of pressure, with an electrical energy consumption of kW (exergy). determine a) the heat flux that must be extracted to produce this ice flux, b) the optimal useful work of the process, c) if this machine is possible or not, d) if the machine is possible, the efficiency of the second law of it. the environment is at 20 °C and 1 bar



1
Expert's answer
2021-10-01T18:12:56-0400

a)

"Q=c_1m(t_1-0)+\lambda m+c_2m(0-t_2)=7.1~\frac Js,"

b)

"\eta=\frac QW=0.035=3.5\%,"

c)

possible

d)

"t=t_1-t_2=24\u00b0C."